# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

### v2

@@ V-B1 Comparing with DeepSpeed-Chat
Interleaving vs Flattening Strategy. We implement two Interleaving strategies to compare them with the Flattening strategy used in DeepSpeed-Chat. In , the Ref and Reward models are placed on half of the GPU devices. In , we interleave not only the Ref and Reward models but also the Actor and Critic models. Based on our experiments as illustrated in Table I, we conclude that the strategy is 12% faster than the Flattening strategy in DeepSpeed-Chat in small training scenarios, such as 7B model. Additionally, the strategy is 71% faster than DeepSpeed-Chat in larger training scenarios, specifically with 65B models when employing ZeRO-3. The speedup ratio further increases when scaling out to more devices.
@@
i) vs. Flattening Strategy. During the training of 7B size models using 18 devices, we observe that strategy in our APP framework leads to a throughput increase from 16.12 to 17.35 samples per second, representing a speedup of 7.63% compared with Flattening strategy in DeepSpeed-Chat. This speedup ratio further improves to 12% when scaling out from 18 to 48 devices. The reason behind this improvement is that the adoption of the interleaving strategy effectively halved the memory redundancy of the Ref and Reward models. Consequently, the freed-up GPU memory can be utilized to expand the overall batch size, as explained in Section IV-A. In this case, the Interleaving strategy achieves a maximum batch size of 36, whereas DeepSpeed-Chat’s maximum batch size is 24. However, it is important to note that does not perform well with larger model scales, such as in the 33B and 65B training scenarios. This is because these models require ZeRO-3 to reduce the memory redundancy to enable training large language models, while there is no memory redundancy of model states in this case.
@@
ii) vs. Flattening Strategy. In contrast, the strategy compensates for this drawback in medium to large training scenarios. Although the strategy only allocates half of the original GPU count for the Actor model, it exhibits significant improvements when combined with the ZeRO-3 mode. As shown in Table I, the strategy is respectively 52% and 71% faster than DeepSpeed-Chat in the 33B and 65B model training scenarios when using 28 devices. The reason behind this improvement is that, under , the Actor and Reward models are placed exclusively on a single node, thus avoiding frequent inter-node AllGather communication primitives in both the prediction and training stages under ZeRO-3. Additionally, the number of communication participants for the Actor/Critic models is reduced by half.
@@
Separation vs Flattening Strategy. The experiments for the Separation strategy are specifically conducted in large-scale training scenarios, such as training 33B and 65B models on and GPU devices. This is because this strategy is not well-suited for small training scenarios with a limited number of GPUs and smaller models. The reason is that the Separation strategy involves adding shadow Actor and Critic models exclusively for the prediction stage, which results in increased memory consumption.
However, the benefits of the Separation strategy are significant, as demonstrated in Table II. Our Separation strategy outperforms the Flattening strategy used in DeepSpeed-Chat by factors ranging from 4 to 11. For instance, in 33B model training using 48 GPU devices, the throughput increases from 0.93 to 4.76 samples per second, resulting in a speedup ratio of 4. The speedup ratio further increases to 7 when the model size is increased to 65B. Additionally, the Separation strategy achieves even higher speedup when scaling out to GPU devices, reaching a speedup ratio of 11 compared to the baseline using DeepSpeed-Chat. The contribution comes from that the Separation strategy decouples the training and inference runtime of the Actor model, which is not possible for the Flattening strategy. This decoupling allows the usage of intra-node tensor parallelism for the generation stages as discussed in section IV-A. For example, in our training of 65B model using GPUs, devices are exclusively allocated for the Shadow Actor model. In this case, we apply tensor parallelism within the same node while replicating it across three machines. This approach significantly reduces inter-node communications compared to frequent inter-node AllGather and Scatter communications in ZeRO-3, which are required in the 33B and 65B model training for the Flattening strategy.
@@
| Separation | 10.07 | 6.80 | |
| # of | Strategy | Model Size |
| GPUs | 33B | |
| DeepSpeed-Chat | 0.69 | |
| Heterogeneous Separation | 0.79 | |
Heterogeneous Separation vs DeepSpeed-Chat. We also evaluate the Heterogeneous Separation strategy in a heterogeneous GPU cluster. The cluster consists of 8 A100 GPUs and an additional 8 spare V100 GPUs with 32GB memory. These heterogeneous V100 GPUs can not be utilized under the Flattening strategy. We specifically conduct experiments using large models, such as the 33B model. Under our heterogeneous Separation strategy, the Ref and Reward models are placed on a separate node with 4 V100 GPUs respectively. We enable the intra-node tensor parallelism for both models and allocate two A100 GPUs to two Shadow Actor replicas exclusively. This configuration aims to accelerate the generation speed and improve the overall training speed.
As shown in Table III, introducing extra V100 devices under the heterogeneous Separation strategy results in a speedup of 14.49%. This acceleration increases the throughput from 0.69 to 0.79 samples per second. Although the acceleration may not be considerable due to the computing power gap between A100 and V100, this strategy explores a novel solution that allows RLHF training to utilize extra heterogeneous resources to enhance the overall throughput. It re-balances the inference workloads among different GPU devices by simply adjusting the batch size, without requiring extensive consideration of software and hardware heterogeneity among different GPUs.
V-B2 Comparing with trlX
In this subsection, we compare our APP framework against trlX, which employs a variant of the Flattening strategy, in the AC-Share scenario. We evaluate our implementation of Interleaving strategies in the APP framework in terms of throughput under various computing resources. However, before presenting the experimental results, there are three important notes to mention:
i) To align the number of models with trlX for a fair comparison, our Interleaving strategy only considers the Actor, Ref, and Reward models, excluding the Critic model in terms of placement. The reason is that the Actor and Critic models share most of the parameters in AC-Share structure. The modifications required are minimal, thanks to the modular design of our model placement abstraction.
ii) We do not include comparative experiments involving the Separation or Heterogeneous Separation strategies because they are primarily effective for models with large parameter sizes, such as the 33B and 65B models. However, trlX can only place the Reward model on a single GPU, requiring the use of offloading techniques[19] to run the 33B and 65B models. However, offloading will significantly reduce the model’s prediction speed. Hence, it can be inferred that our Separation strategy offers even greater advantages in the trlX scenario compared to experiment results in DeepSpeed-Chat.
@@ V-B2 Comparing with trlX
Specifically, our Interleaving strategy can achieve a speedup ratio of up to compared with the trlX, especially with gradient checkpointing enabled 444Gradient checkpointing[4] refers to whether activation value recalculations are enabled during training. and more participating nodes.
Comparison Under Gradient Checkpointing. As shown in Table IV, our Interleaving strategy without gradient checkpointing accelerated the throughput from 6.04 to 11.17 samples per second, resulting in a speedup of 84.93% for training the 6B models. This improvement arises by avoiding the single-device bottleneck in trlX, where one device is responsible for aggregating data from other devices, incurring significant communication costs and leaving other devices idle, as discussed in Section III-A. Moreover, the remaining memory that is not utilized by the Reward model is wasted, limiting the overall batch size. Also, this design leads to a longer generation duration, leaving other devices idle and resulting in underutilized resources, regardless of whether trlX is in standalone or coexisted mode. It is noteworthy that trlX narrows the performance gap when the Reward model can fully utilize the single device in 13B model training. Additionally, when gradient checkpointing is enabled, the performance gap increases significantly, resulting in a speedup ratio of 360.64%. This is because gradient checkpointing can reduce memory consumption and in turn, increase the overall training batch size. However, the single-device bottleneck in trlX disables this benefit, hindering the potential performance improvement from gradient checkpointing.
Comparison When Scaling Out. Our Interleaving strategy demonstrates significant improvements in training speed compared to trlX, which achieves acceleration ranging from to , when scaling out. As shown in Table V, our Interleaving strategy increases the throughput from 6.21 to 27.27 samples per second, resulting in a speedup ratio of 3.4 compared to trlX for the 7B model training on 18 GPUs. This speedup ratio further increases to 6.1 and 11 when scaling out to 28 and 48 GPU devices respectively. The training with the 13B models exhibits a similar pattern. It is analyzed that the single-GPU bottleneck in the trlX approach significantly limited the throughput. Furthermore, the interleaving of the Reward model and Ref model in our Interleaving strategy reduces memory redundancy as discussed in Section IV-A, resulting in considerable acceleration.
@@ V-B3 Detailed Analysis
i) In the AC-NonShare scenario, we experiment with the Llama 33B model on 48 GPU devices. Under DeepSpeed-chat’s Flattening strategy, the generation stage occupies the majority of the total duration, accounting for 93.31% in one step, while the training stage only takes up 6.69% as illustrated in Figure 9. However, our Separation strategy reduces the overall duration in one step from around 400 seconds to 30 seconds by significantly reducing the time-consuming generation stage. The percentage of the generation stage is reduced from 93.31% to 56%, almost half of the total duration. This reduction is attributed to our Separation strategy, which allows for intra-node tensor parallelism and other inference-optimized techniques for the inference models. Additionally, our Interleaving strategy increases 50% of the maximum batch size available, from 24 to 36 compared with the DeepSpeed-Chat in the 7B model, as displayed in Figure 9.
ii) In the AC-Share scenario, we examine the time expenses for the Llama 7B model on 28 GPU devices. As depicted in Figure 9, our approach can reduce the overall duration from around 72 seconds to 10 seconds by largely decreasing the generation stage, as the percentage of generation duration drops from 97.86% to 34.83%. In the generation phase, the trlX approach suffers from the aforementioned single-device bottleneck. In contrast, our Interleaving strategy can speed up the generation stage by placing the Reward model on multiple devices and interleaving it with the Ref model. For the maximum batch size available, the Interleaving strategy increases the batch size 25 compared with trlX in the 13B model, rocketing up from 2 to 50, as illustrated in Figure 9.
@@ VI Conclusion
In this paper, we introduce a novel RLHF training framework that adaptively places multiple large language models in the RLHF pipeline according to different training scenarios in terms of model and device scales. Our two placement strategies, Interleaving and Separation, not only significantly enhance training efficiency but also offer adaptable solutions for different training scenarios. Additionally, our framework is easy to use and provides simple user interface.

### v3

@@ V-B1 Comparison in AC-NonShare Structure
We implement two Interleaving strategies to compare them with the Co-located strategy used in DeepSpeed-Chat. In , the Ref and Reward models are placed on half of the GPU devices. In , we interleave not only the Ref and Reward models but also the Actor and Critic models. Based on our experiments as illustrated in Table I, we conclude that the strategy is 12% faster than the Co-located strategy in DeepSpeed-Chat in small training scenarios, such as 7B model. Additionally, the strategy is 71% faster than DeepSpeed-Chat in larger training scenarios, specifically with 65B models when employing ZeRO-3. The speedup ratio further increases when scaling out to more devices.
@@
i) vs. Co-located Strategy. During the training of 7B size models using 18 devices, we observe that strategy in our FlexRLHF framework accelerates the throughput from 16.12 to 17.35 samples per second, representing a speedup of 7.63% compared with Co-located strategy in DeepSpeed-Chat. This speedup ratio further improves to 12% when scaling out from 18 to 48 devices. The reason behind this improvement is that the adoption of the interleaving strategy effectively halved the memory redundancy of the Ref and Reward models. Consequently, the freed-up GPU memory can be utilized to expand the overall batch size, as explained in Section IV-A. In this case, the Interleaving strategy achieves a maximum batch size of 36, whereas DeepSpeed-Chat’s maximum batch size is 24. However, it is important to note that does not perform well with larger model scales, such as in the 33B and 65B training scenarios. This is because these models require ZeRO-3 to reduce the memory redundancy to enable training large language models, while there is no memory redundancy of model states in this case.
@@
ii) vs. Co-located Strategy. In contrast, the strategy compensates for this drawback in medium to large training scenarios. Although the strategy only allocates half of the original GPU count for the Actor model, it exhibits significant improvements when combined with the ZeRO-3 mode. As shown in Table I, the strategy is respectively 52% and 71% faster than DeepSpeed-Chat in the 33B and 65B model training scenarios when using 28 devices. The reason behind this improvement is that, under , the Actor and Reward models are placed exclusively on a single node, thus avoiding frequent inter-node AllGather communication primitives in both the prediction and training stages under ZeRO-3. Additionally, the number of communication participants for the Actor/Critic models is reduced by half.
@@
Disaggregated vs Co-located Strategy. The experiments for the Disaggregated strategy are specifically conducted in large-scale training scenarios, such as training 33B and 65B models on and GPU devices. This is because this strategy is not well-suited for small training scenarios with a limited number of GPUs and smaller models. The reason is that the Disaggregated strategy involves adding shadow Actor and Critic models exclusively for the prediction stage, which results in increased memory consumption.
However, the benefits of the Disaggregated strategy are significant, as demonstrated in Table II. Our Disaggregated strategy outperforms the Co-located strategy used in DeepSpeed-Chat by factors ranging from 4 to 11. For instance, in 33B model training using 48 GPU devices, the throughput increases from 0.93 to 4.76 samples per second, resulting in a speedup ratio of 4. The speedup ratio further increases to 7 when the model size is increased to 65B. Additionally, the Disaggregated strategy achieves even higher speedup when scaling out to GPU devices, reaching a speedup ratio of 11 compared to the baseline using DeepSpeed-Chat. The contribution comes from that the Disaggregated strategy decouples the training and inference runtime of the Actor model, which is not possible for the Co-located strategy. This decoupling allows the usage of intra-node tensor parallelism for the generation stages as discussed in section IV-A. For example, in our training of 65B model using GPUs, devices are exclusively allocated for the Shadow Actor model. In this case, we apply tensor parallelism within the same node while replicating it across three machines in a DP manner. This approach significantly reduces inter-node communications compared to frequent inter-node AllGather and Scatter communications in ZeRO-3 and resharding in Hybrid Engine.
@@
| Disaggregated | 10.07 | 6.80 | |
| # of | Strategy | Model Size | |
| GPUs | 33B | ||
| DeepSpeed-Chat | 0.69 | ||
| Heterogeneous Disaggregated | 0.79 | ||
Heterogeneous Disaggregated vs Co-located Strategy.
We also evaluate the Heterogeneous Disaggregated strategy in a heterogeneous GPU cluster. The cluster consists of 8 A100 GPUs and an additional 8 spare V100 GPUs with 32GB memory. These heterogeneous V100 GPUs can not be easily utilized under the Co-located strategy. We specifically conduct experiments using large models, such as the 33B model. Under our heterogeneous Disaggregated strategy, the Ref and Reward models are placed on a separate node with 4 V100 GPUs respectively. We enable the intra-node tensor parallelism for both models and allocate two A100 GPUs to two Shadow Actor replicas exclusively. This configuration aims to accelerate the generation speed and improve the overall training speed.
As shown in Table III, introducing extra V100 devices under the heterogeneous Disaggregated strategy results in a speedup of 14.49%. This acceleration increases the throughput from 0.69 to 0.79 samples per second. Although the acceleration may not be considerable due to the computing power gap between A100 and V100, this strategy explores a novel solution that allows RLHF training to utilize extra heterogeneous resources to enhance the overall throughput. It re-balances the inference workloads among different devices by simply adjusting the batch size, without consideration of software and hardware heterogeneity among different GPUs.
V-B2 Comparison in AC-Share Structure
In the AC-Share scenario, we compare our FlexRLHF framework against trlX. We evaluate our implementation of Interleaving strategies in the FlexRLHF framework in terms of throughput under various computing resources. However, before presenting the experimental results, there are three important notes to mention:
i) To align the number of models with trlX for a fair comparison, our Interleaving strategy only considers the Actor, Ref, and Reward models, excluding the Critic model in terms of placement. The reason is that the Actor and Critic models share most of the parameters in AC-Share structure. The modifications required are minimal, with help of the modular design of our model placement abstraction.
ii) We do not include comparative experiments involving the Disaggregated or Heterogeneous Disaggregated strategies because they are primarily effective for models with large parameter sizes, such as the 33B and 65B models. However, trlX can only place the Reward model on a single GPU, requiring the use of offloading techniques [25] to run the 33B and 65B models. However, offloading will significantly reduce the model’s prediction speed. Hence, it can be inferred that our Disaggregated strategy offers more advantages in the trlX scenario compared to results in DeepSpeed-Chat.
@@ V-B2 Comparison in AC-Share Structure
Comparison under Gradient Checkpointing. As shown in Table IV, our Interleaving strategy without gradient checkpointing accelerated the throughput from 6.04 to 11.17 samples per second, resulting in a speedup of 84.93% for training the 6B models. This improvement arises by avoiding the single-device bottleneck of the Reward model in trlX, as discussed in Section III-B. Moreover, the remaining memory that is not utilized by the Reward model is wasted, limiting the overall batch size. Also, this design leads to a longer generation duration, leaving other devices idle and resulting in underutilized resources, regardless of whether trlX is in standalone or coexisted mode. It is noteworthy that trlX narrows the performance gap when the Reward model can fully utilize the single device in 13B model training. Additionally, when gradient checkpointing is enabled, the performance gap increases significantly, resulting in a speedup ratio of 360.64%. This is because gradient checkpointing can reduce memory consumption and in turn, increase the overall training batch size. However, the single-device bottleneck in trlX disables this benefit, hindering the potential performance improvement from gradient checkpointing.
Scalability. Our Interleaving strategy demonstrates significant improvements in training speed compared to trlX, which achieves acceleration ranging from to , when scaling out. As shown in Table V, our Interleaving strategy increases the throughput from 6.21 to 27.27 samples per second, resulting in a speedup ratio of 3.4 compared to trlX for the 7B model training on 18 GPUs. This speedup ratio further increases to 6.1 and 11 when scaling out to 28 and 48 GPU devices respectively. The training with the 13B models exhibits a similar pattern. It is analyzed that the single-GPU bottleneck in the trlX approach significantly limited the throughput. Furthermore, the interleaving of the Reward model and Ref model in our Interleaving strategy reduces memory redundancy as discussed in Section IV-A, resulting in considerable acceleration.
@@ V-B3 Detailed Analysis
i) In the AC-NonShare scenario, we experiment with the Llama 33B model on 48 GPU devices. Under DeepSpeed-chat’s Co-located strategy, the generation stage occupies the majority of the total duration, accounting for 93.31% in one step, while the training stage only takes up 6.69% as illustrated in Fig. 9. However, our Disaggregated strategy reduces the overall duration in one step from around 400 seconds to 30 seconds by significantly reducing the time-consuming generation stage. The percentage of the generation stage is reduced from 93.31% to 56%, almost half of the total duration. This reduction is attributed to our Disaggregated strategy, which allows for intra-node tensor parallelism and other inference-optimized techniques for the inference models without AllGather parameters for each generation. Additionally, our Interleaving strategy increases 50% of the maximum batch size available, from 24 to 36 compared with the DeepSpeed-Chat in the 7B model, as displayed in Fig. 9.
ii) In the AC-Share scenario, we examine the time expenses for the Llama 7B model on 28 GPU devices. As depicted in Fig. 9, our approach can reduce the overall duration from around 72 seconds to 10 seconds by largely decreasing the generation stage, as the percentage of generation duration drops from 97.86% to 34.83%. In the generation phase, the trlX approach suffers from the aforementioned single-device bottleneck. In contrast, our Interleaving strategy can speed up the generation stage by placing the Reward model on multiple devices and interleaving it with the Ref model. For the maximum batch size available, the Interleaving strategy increases the batch size 25 compared with trlX in the 13B model, rocketing up from 2 to 50, as illustrated in Fig. 9.
@@ VI Conclusion
In this paper, we introduce a novel RLHF training framework that flexibly places multiple LLMs according to the characteristics of RLHF in different training scenarios in terms of model size and device scales. Our two placement strategies, Interleaving and Disaggregated strategies, not only significantly improve training efficiency but also offer agile solutions for various training scenarios. Additionally, our framework is easy to use and provides a simple user interface in practice.